OpenStudy (anonymous):

need help simplifying -15xy^2/20x^6y^6

ganeshie8 (ganeshie8):

\(\huge \frac{-15xy^2}{20x^6y^6}\)

ganeshie8 (ganeshie8):

first cancel the common factors in number constants... if any..

ganeshie8 (ganeshie8):

5 goes in both 15 and 20 right ?

OpenStudy (anonymous):

yes so i have \[-3/4x^5y^5\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i mean y^4

ganeshie8 (ganeshie8):

\(\huge \frac{-15xy^2}{20x^6y^6}\) \(\huge \frac{-3xy^2}{4x^6y^6}\) \(\huge \frac{-3}{4x^5y^4}\)

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

np :)